OpenStudy (anonymous):

given that the x-intercept of a line is (1,0) and y-intercept is (0,-3), what is the equation of the line in slope-intercept form

OpenStudy (the_fizicx99):

Well, first find the slope. M = y2-y1/x2-x1 Then plug in: y-y1=M(x-x1) and solve for that, it'll give you the slope intercept form equation.

OpenStudy (anonymous):

Thats really confusing to me

OpenStudy (the_fizicx99):

Just plug in the value, -3-0/0-1 -3 - 0 = -3 0 - 1 = -1 So -3/-1 = 3 <-- slope.. M Now plug in the values for: y-y1=M(x-x1)

OpenStudy (anonymous):

so would it be like y--3=3(x-0) than just work it out

OpenStudy (the_fizicx99):

You got your values mixed up.. its (x,y) y-0=3(x-1)

OpenStudy (anonymous):

so would my answer be y=3x-1?

OpenStudy (the_fizicx99):

No.. what is the 3 doing to (x-1)??? Distributive property so you multiply 3*x and 3* (-1)

OpenStudy (anonymous):

so y=3x-3?

OpenStudy (the_fizicx99):

yes

OpenStudy (anonymous):

Thank you.!

OpenStudy (the_fizicx99):

Yw